# Hyperuniformity and optimal transport of point processes

### v3

@@ 3.4 Proofs of Theorems 1 and 2
Consider first the case $d\geq 3$ with the weight function $w(x)=x^{2}$. By Theorem 3, we have that there exists a transport plan $m_{n}$ such that $\mathbf{E}\widetilde{{\mathsf{C}}}_{w}(\tilde{\mu}_{n},\tilde{\mathcal{L}}_{n};\Lambda_{n})\leqslant cn$ for all $n\geqslant 1$ and similarly for $\tilde{\nu}_{n}$.
Now using Proposition 11, there exist point processes $\mu^{\prime},\nu^{\prime}$ on probability spaces $(\Omega,\mathbf{P})$, $(\Omega^{\prime},\mathbf{P}^{\prime})$ respectively and matchings $T:\mathbb{Z}^{d}+U\to{\rm supp}(\mu^{\prime}),T^{\prime}:\mathbb{Z}^{d}+U^{\prime}\to{\rm supp}(\nu^{\prime}),$ with $U,U^{\prime}$ being uniformly distributed in $[0,1]^{d}$ respectively on $\Omega,\Omega^{\prime}$, such that the corresponding typical matching distances $X,X^{\prime}$ satisfy $\mathbf{E}w(X)<\infty,\mathbf{E}^{{}^{\prime}}w(X^{\prime})<\infty$. On the product space $\Omega\times\Omega^{\prime}$ endowed with $\mathbf{P}\otimes\mathbf{P}^{\prime}$, we have an obvious invariant matching $S:\mathbb{Z}^{d}+U\to\mathbb{Z}^{d}+U^{\prime}$ with bounded typical distance, therefore $T^{-1}\circ S\circ T^{\prime}$ is an invariant matching between $\mu^{\prime}$ and $\nu^{\prime}$ whose typical distance $\tilde{X}$ satisfies $\mathbf{E}w(\tilde{X})<\infty$.
@@
As explained in Remark 1, the proof for $d=1,2$ follows from this bound and boundedness of $\sigma_{\mu},\sigma_{\nu}$, which are implied by the integrability of the RPCM $\beta$. As argued in the case of $d=3$ above, this requirement can be further reduced to exhibiting a $\mathbb{Z}^{d}$-translation invariant matching $T$ from $\mu$ to $\nu$ satisfying (44).
The construction in [30, Theorem 1(ii)] proceeds as follows. Let $Z_{1},\ldots,Z_{k},\ldots$ be i.i.d. uniform random vectors in $\{0,1\}^{d}$ independent of $\mu,\nu$. Successively partition $\mathbb{R}^{d}$ into randomly centered (at $\sum_{i=0}^{k-1}2^{i}Z_{i}$) cubes of side-length $2^{k}$ for $k\geq 0$ and call these as cubes at level $k$ i.e., $k$-level cubes are
$2^{k}z+[0,2^{k}]^{d}+\sum_{i=0}^{k-1}2^{i}Z_{i},\,z\in\mathbb{Z}^{d}.$
At level $k=0$ all points are considered unmatched. At any level $k$, as many unmatched points of $\mu$ and $\nu$ within a $k$-level cube are matched as possible. This (partial) matching is carried using an arbitrary rule. The unmatched points after this are carried forward to the next level and the procedure is repeated. The (partial) matching thus induced is said to be $T$ and with the unmatched points, matched to $\infty$. If $T$ satisfies the tail bound (44), then letting $r\to\infty$, we can see that $T$ matches every $\mu$ point to a $\nu$ point and by stationarity and that $\mu,\nu$ have equal intensities, one can argue vice-versa as well; see [30, Lemma 7]. Thus, we have that $T$ is a (perfect) matching between $\mu$ and $\nu$ and by construction $\mathbb{Z}^{d}$-invariant.
Now we show the tail bound (44) for the partial matching $T$ as described above. Using mass transport principle and that the random centering is independent of $\mu,\nu$, one can derive that (see derivation of equation (11) in [30]),
$\mathbf{E}\Big{(}\nu\{x\in[0,1]^{d}:\|T(x)-x\|\geqslant\sqrt{d}2^{k}\}\Big{)}\leqslant 2^{-dk}\mathsf{E}S_{+},$
where $S=(\nu(B)-\mu(B))$ for $B=[0,2^{k})^{d}$ and recall that $S_{+}=\max\{S,0\}$. Since $\mu,\nu$ have equal intensities, we can bound this via Cauchy-Schwarz and triangle inequality as follows.
@@
where recall that $\sigma_{\mu}(B):=|B|^{-1}\text{Var}(\mu(B))$ and similarly $\sigma_{\nu}(B)$. This gives the bound in (44) for $r=\sqrt{d}2^{k}$ and one can then extrapolate it to all $r$. ∎
@@
(1).   Under same assumptions for $d=2$, Theorem 4(2) yields that $\mathbf{E}\widetilde{W}_{2}^{2}(\tilde{\mu}_{n},\tilde{\mathcal{L}}_{n};\Lambda_{n})\leqslant cn$ for all $n\geqslant 1$. Now using Proposition 11 as in the proof of Theorem 1, the conclusion follows. The claims about matching with a general point process $\nu$ has already been justified below the theorem statement; see also the beginning of the proof of Theorem 1.
(2).   For $d=1$ and under the assumption of $|x||\beta(dx)|$ being integrable, we have that $\sigma_{\mu}(r)\leq cr^{-1}$ for some constant $c$; see [45, Proposition 2]. The same bound holds for $\nu$ as well. Now using (44), we obtain that there exists a matching $T$ such that the typical matching distance $X$ satisfies
@@
Now following the arguments as in Remark 1, we obtain the claim as required.

### v4

@@ 3.4 Proofs of Theorems 1 and 2
In both the proofs below, we first justify the case of independent simple point processes $\mu,\nu$ with integrable RPCM and then address the case of shifted lattices. For the latter, we use Example 1, a perturbed lattice that has an integrable RPCM, and allows to bridge the results between the two categories. However, it is possible with some more effort to use [25] as in the last part of proof of Proposition 11 to give a more direct argument.
@@
Consider first the case $d\geq 3$ with the weight function $w(x)=x^{2}$. By Theorem 3, we have that there exists a finite constant $c$ such that $\mathbf{E}\widetilde{{\mathsf{C}}}_{w}(\tilde{\mu}_{n},\tilde{\mathcal{L}}_{n};\Lambda_{n})\leqslant cn$ for all $n\geqslant 1$ and similarly for $\tilde{\nu}_{n}$. By triangle inequality, we have that for some finite constant $c$,
$\mathbf{E}\widetilde{{\mathsf{C}}}_{w}(\tilde{\mu}_{n},\tilde{\nu}_{n};\Lambda_{n})\leq 4\Big(\mathbf{E}\widetilde{{\mathsf{C}}}_{w}(\tilde{\mu}_{n},\tilde{\mathcal{L}}_{n};\Lambda_{n})+\mathbf{E}\widetilde{{\mathsf{C}}}_{w}(\tilde{\nu}_{n},\tilde{\mathcal{L}}_{n};\Lambda_{n})\Big)\leq cn,$
for all $n\geq 1$. Now using Proposition 11, there exist point processes $\mu^{\prime}(\stackrel{{\scriptstyle(d)}}{{=}}\mu),\nu^{\prime}(\stackrel{{\scriptstyle(d)}}{{=}}\nu)$ on the same probability space and a translation-invariant matching $T$ between $\mu^{\prime}$ and $\nu^{\prime}$ such that its typical distance $X$ satisfies $\mathbf{E}w(X)<\infty$.
@@
The proof for $d=1,2$ follows from this bound and boundedness of $\sigma_{\mu},\sigma_{\nu}$, which are implied by the integrability of the RPCM $\beta$. To check this, verify by contradiction that for $x,r>10,x^{d/2}\ln(x)^{-\gamma}>r$ implies $x^{d/2}>cr\ln(r)^{\gamma}$. Now for $\gamma>1$, substituting the above tail probability bound (50) along with boundedness of $\sigma_{\mu},\sigma_{\nu}$ yields that
$\mathbf{E}\left(\frac{X^{d/2}}{1+|\ln X|^{\gamma}}\right)$ $\leqslant 10+\int_{10}^{\infty}\mathbf{P}\Big(\frac{X^{d/2}}{1+|\ln X|^{\gamma}}\geq r\Big)\mathrm{d}r$
$\leqslant 10+\int_{10}^{\infty}\mathbf{P}(X^{d/2}\geqslant cr\ln r^{\gamma})\mathrm{d}r<\infty$
$\leqslant 10+\int_{10}^{\infty}\frac{c}{r(1+(\ln r)^{\gamma})}\mathrm{d}r<\infty,$
which concludes the proof. The requirement of (9) can be further reduced to exhibiting a $\mathbb{Z}^{d}$-translation invariant matching $T$ from $\nu$ to $\mu$ satisfying (50). Thanks to (2), (50) is implied by
$\mathbf{E}\Big(\nu\{x\in[0,1]^{d}:\|T(x)-x\|\geqslant r\}\Big)\leqslant cr^{-d/2}\sqrt{\sigma_{\mu}(r)+\sigma_{\nu}(r)},\,\,r\geq 1.$ (51)
The construction in [33, Theorem 1(ii)] proceeds as follows. Let $Z_{k},k\geqslant 0$ be i.i.d. uniform random vectors in $\{0,1\}^{d}$ independent of $\mu,\nu$. We successively partition $\mathbb{R}^{d}$ into randomly centered cubes of side-length $2^{k}$ and randomly match points iteratively. More precisely, let ${U_{k}}=\sum_{i=0}^{k-1}2^{i}Z_{i}$ for $k\geq 0$ and define the $k$-level cubes as the
$2^{k}z+U_{k}+[0,2^{k}]^{d},z\in\mathbb{Z}^{d}.$
At level $k=0$ all points are considered unmatched. At any level $k\geqslant 1$, for any $k$-level cube $C$, let $m_{C,\mu},m_{C,\nu}$ the number of yet unmatched points of $\mu\cap C,\nu\cap C$ respectively. Form $m_{C}:=\min(m_{C,\mu},m_{C,\nu})$ pairs of matched points of $\mu$ and $\nu$ within $C$ using an arbitrary rule ; for example, we can use the stable marriage algorithm as in [31]. The points still unmatched after this step are carried forward to the next level $k+1$ and the procedure is repeated. Eventually, all points of either $\nu$ or $\mu$ are matched, all remaining points (i.e., the unmatched points) are said to be matched to $\infty$. Denote by $T$ the corresponding matching from $\nu\cup\{\infty\}$ to $\mu\cup\{\infty\}$. If $T$ satisfies the tail bound (51), then letting $r\to\infty$, we can see that a.s. every point is matched to a finite distance, i.e. $T$ matches every $\nu$ point to a $\mu$ point, one can argue vice-versa as well; see [33, Lemma 7]. Thus, we have that $T$ is a (perfect) matching between $\mu$ and $\nu$ and by construction $\mathbb{Z}^{d}$-invariant.
Now we show the tail bound (51) for the partial matching $T$ as described above. The proof relies on the estimate
$\mathbf{E}\Big({\nu}\{x\in[0,1]^{d}:\|T(x)-x\|\geqslant\sqrt{d}2^{k}\}\Big)\leqslant 2^{-dk}\mathsf{E}S_{+}$
where $S=(\nu(B)-\mu(B))$ for $B=[0,2^{k})^{d}$ and recall that $S_{+}=\max\{S,0\}$. We refer the reader to the geometric proof of equation (11) in [33] as it is exactly identical. Note that $\sqrt{d}2^{k}$ bounds the diameter of any $k$-level cube. In the above derivation, a crucial observation is that for $x\in\nu$ to be unmatched or matched to a $\mu$-point at distance larger than $\sqrt{d}2^{k}$, it cannot be matched within the $k$-level cube $C$ containing $x$. This can happen only if there are excess $\nu$ points than $\mu$ in the $k$-level cube $C$. Hence $x\in\nu$ has matching distance larger than $\sqrt{d}2^{k}$ only if $\nu(C)>\mu(C)$. Then relying on the mass transport principle and the fact that the random centering is independent of $\mu,\nu$, one can derive the above bound.
Since $\mu,\nu$ have equal intensities, we can bound this via Cauchy-Schwarz and triangle inequality as follows.
@@
where recall that $\sigma_{\mu}(B):=|B|^{-1}\text{Var}(\mu(B))$ and similarly $\sigma_{\nu}(B)$. This gives the bound in (51) for $r=\sqrt{d}2^{k}$ and one can then extrapolate it to all $r$.
Now we justify the case of the shifted lattice $\mathbb{Z}^{d}+U$ with $U$ being uniformly distributed in $0,1]^{d}$. Let $\nu_{\textrm{Cl}}$ be the cloaked stationary i.i.d. perturbed lattice in $d\geq 1$ as defined in Example 1. Note that it has an integrable RPCM. We have proven that there exists a translation-invariant matching $T_{1}$ between $\mu$ and $\nu_{\textrm{Cl}}$ (independently constructed on a probability space) such that its typical distance $X$ satisfies $\mathbf{E}w(X_{1})<\infty$. By construction of $\nu_{\textrm{Cl}}$, there exists a translation-invariant matching $T_{2}$ between $\mathbb{Z}^{d}+U$ and $\nu_{\textrm{Cl}}$ such that its typical distance $X_{2}$ satisfies $X_{2}\leq 2\sqrt{d}$ a.s.. Thus $T=T_{1}\circ T_{2}$ gives a translation-invariant matching between $\mathbb{Z}^{d}+U$ to $\mu$ and using triangle inequality as above, we have that the typical distance $X$ satisfies $\mathbf{E}w(X)<\infty$. ∎
@@
(1). Since our assumption on point processes $\mu,\nu$ for $d=2$ is exactly the same as that in Theorem 4(2), we obtain that $\mathbf{E}\widetilde{W}_{2}^{2}(\tilde{\mu}_{n},\tilde{\mathcal{L}}_{n};\Lambda_{n})\leqslant cn$ for all $n\geqslant 1$ and similarly for $\tilde{\nu}_{n}$. Now via triangle inequality as in Theorem 1, we derive that for some finite constant $c$,
$\mathbf{E}\widetilde{W}_{2}^{2}(\tilde{\mu}_{n},\tilde{\nu}_{n};\Lambda_{n})\leqslant cn,\,\forall n\geq 1.$
Now using Proposition 11, we obtain that there exist point processes $\mu^{\prime}(\stackrel{{\scriptstyle(d)}}{{=}}\mu),\nu^{\prime}(\stackrel{{\scriptstyle(d)}}{{=}}\nu)$ on the same probability space and a translation-invariant matching $T$ between $\mu^{\prime}$ and $\nu^{\prime}$ whose typical distance $X$ satisfies $\mathbf{E}X^{2}<\infty$. Thus the conclusion follows for independent HU point processes $\mu,\nu$ with integrable RPCM and satisfying the logarithmic integrability condition..
The case of shifted lattice $\mathbb{Z}^{d}+U$ can again be justified as at the end of the proof of Theorem 1. In particular, the stationary i.i.d. perturbed lattice $\nu_{\textrm{Cl}}$ in Example 1 is hyperuniform and its RPCM satisfies the logarithmic integrability condition as well since it is compactly supported. Thus there exists a translation-invariant matching $T_{1}$ between $\mu$ and $\nu_{\textrm{Cl}}$ whose typical distance $X_{1}$ satisfies $\mathbf{E}X_{1}^{2}<\infty$. Now using the canonical matching $T_{2}$ between $\mathbb{Z}^{d}+U$ and $\nu_{\textrm{Cl}}$, we can justify existence of a translation-invariant matching $T:\mathbb{Z}^{d}+U\to\mu$ with typical distance $X$ satisfying $\mathbf{E}X^{2}<\infty$.
(2).   For $d=1$. Assume that the number variance is bounded for $\mu$ and $\nu$. Now using (50), we obtain that there exists a translation-invariant matching $T$ such that the typical matching distance $X$ satisfies
@@
Notice that for $r>1$, if $X/(1+|\ln X|^{\gamma})\geqslant r$, then $X\geqslant r$ and so $X\geqslant r(1+\ln r^{\gamma})$. Now for $\gamma>1$, substituting the above tail probability bound yields that
$\mathbf{E}\left(\frac{X}{1+|\ln X|^{\gamma}}\right)$ $\leqslant 1+\int_{1}^{\infty}\mathbf{P}\Big(\frac{X}{1+|\ln X|^{\gamma}}\geq r\Big)\mathrm{d}r$
$\leqslant 1+\int_{1}^{\infty}\mathbf{P}(X\geqslant r(1+(\ln r)^{\gamma}))\mathrm{d}r<\infty$
$\leqslant 1+\int_{1}^{\infty}\frac{c}{r(1+(\ln r)^{\gamma})}\mathrm{d}r<\infty,$
which concludes the proof. ∎